Circular No. 6363

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) Contrary to what was implied, the analysis in the second paragraph of

IAUC 6360 by Z. Sekanina of a highly decelerated fragment of the comet’s nucleus does not involve the observations discussed in the first paragraph. With reference to those observations, and amplifying the remark about the Hubble Space Telescope observations, H. A. Weaver, Applied Research Corporation, reports the following, on behalf of his HST team: “A detailed analysis of several R-band HST WFC images of the comet taken near Mar. 26.063 UT shows no evidence for companion condensations of the type described in the first paragraph of IAUC 6360, either in the raw images or after division by a symmetrical model coma. WFC pixels are 0′′.0996 across, and this projects to a distance of 7.6 km at the comet. Multiple nuclei with separations of > ∼ 0′′.2 would be easily resolved in the HST images. The inner coma is reminiscent of the coma of 1P/Halley during the Giotto encounter, with multiple jets in the sunward-facing hemisphere, the strongest jet pointing directly toward the projected solar vector. In contrast to 1P/Halley there is in C/1996 B2 a bright jet that appears to emanate from the night side of the nucleus, but we cannot rule out the possibility that the emission is really on the illuminated portion and only appears to be on the night side because of projection effects (the solar phase angle being 72^{o}). The remnants of the “knots” in the tail detected two days earlier by Lecacheux et al. (IAUC 6354) are clearly seen in the HST images. We identify at least three separate entities: a “streak” having a sharp sunward-pointing spike, a faint condensed feature and an elongated “puff”. The offsets from the nucleus are Δα = −8′′.6, Δδ = −14′′.6; −8′′.4, −16′′.6; and −9′′.4,−18′′.8, respectively. We cannot yet say anything definite regarding the size of the nucleus.”

H. E. Matthews, Herzberg Institute of Astrophysics, Ottawa, and Joint Astronomy Centre, Hilo (JAC); D. Jewitt, M. Senay and T. Owen, University of Hawaii; J. K. Davies and W. R. F. Dent (JAC); and N. Biver, D. Bockelée-Morvan, J. Crovisier, H. Rauer and D. Gautier, Observatoire de Paris-Meudon, communicate: “We detected submillimeter continuum emission from C/1996 B2 using the James Clerk Maxwell Telescope on Mauna Kea. Preliminary estimates of the flux density at 350 and 800 μm are 6.8 ± 1.3 Jy and 0.55 ± 0.04 Jy, respectively, on both Mar. 23 and 24 UT. These are by far the strongest continuum detections ever achieved in a comet. From these and other measurements we deduce a a relatively flat spectral index, −2.9 ± 0.1. The estimated dust mass in the 19′′ (1700 km) diameter beam is of order 10^{8} kg. Assuming a dust outflow velocity of 10 m/s, the dust production rate is 1200 kg/s.”

1996 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden